adrienne invested a total amount of 100,000 php in banks A and B with an interest rate of 5% and 8% respectively. if the income from bank B is 800 more than the twice the income from bank A, how much was invested in bank B?
Found 3 solutions by Boreal, MathTherapy, ikleyn:
Answer by Boreal(15235)

x=5% amount
100000-x=8% amount
.05x*2=(0.08)(100000-x), since B is twice as large as A
.10x=-0.08x+8000
.18x=8000
x=44,444.44 (*0.05=2222.22)
100000-x=55,555.56 (*.08=4444.44)
55,555.56 ANSWER

Answer by MathTherapy(10552)


adrienne invested a total amount of 100,000 php in banks A and B with an interest rate of 5% and 8% respectively. if the income from bank B is 800 more than the twice the income from bank A, how much was invested in bank B?
That other person is WRONG, WRONG, WRONG, WRONG, WRONG, WRONG, WRONG, and WRONG, again!!!
Correct answer:
ACCEPT NO OTHER!!!

Answer by ikleyn(52781)

.
x = 5% amount
100000-x = 8% amount


0.08*(100000-x) = 0.05*x*2+800,     since it is given


8000 - 800 = 0.08x + 0.10x

0.18x = 7200

x = 7200/0.18 = 40000  was invested at 5%.
  
The rest, i.e. 100000-40000 = 60000 was invested at 8%.
  

ANSWER.  Php 40000 was invested at 5%;  Php 60000 was invested at 8%.


Check.   0.08*60000 - 2*(0.05*40000) = 800 Php   ! Correct !

Solved.

-----------------

It is a standard and typical problem on investments.

If you need more details,  or if you want to see other similar problems solved by different methods,  look into the lesson
    - Using systems of equations to solve problems on investment
in this site.

You will find there different approaches  (using one equation or a system of two equations in two unknowns),  as well as
different methods of solution to the equations  (Substitution,  Elimination).

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic  "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.